# Towards Fixing Panic Bugs for Real-world Rust Programs

### v1

@@ 5.2. Answer to RQ2: Fixing Effectiveness
Table 7. The number of correct patches and correctness of different tools on Panic4R with perfect locations, respectively.
@@
Plausible Correct Plausible Correct Plausible Correct
Top-1 GPT-4 6 9.8% 4 6.6% 3 7.3% 3 7.3% 9 8.8% 7 6.9%
GPT4-multi 6 9.8% 4 6.6% 3 7.3% 3 7.3% 9 8.8% 7 6.9%
PanicKiller 33 54.1% 24 39.3% 10 24.4% 7 17.1% 43 42.2% 31 30.4%
Top-3 GPT-4 6 9.8% 5 8.2% 3 7.3% 3 7.3% 9 8.8% 8 7.8%
GPT4-multi 6 9.8% 5 8.2% 3 7.3% 3 7.3% 9 8.8% 8 7.8%
PanicKiller 36 59.0% 27 44.3% 13 31.7% 11 26.8% 49 48.0% 38 37.3%
Top-5 GPT-4 6 9.8% 5 8.2% 4 9.8% 4 9.8% 10 9.8% 9 8.8%
GPT4-multi 6 9.8% 5 8.2% 4 9.8% 4 9.8% 10 9.8% 9 8.8%
PanicKiller 36 59.0% 27 44.3% 13 31.7% 11 26.8% 49 48.0% 38 37.3%
Regarding to the correct and plausible patches, we believe that, when coupled with the textual interpretations derived from our mined patterns, plausible patches can aid developers repair their program efficiently. This is because the failure of some patches to pass the regression tests may be attributed to additional logic after the patch is applied, which is beyond APR’s capabilities. Such plausible fixes, with minimal manual oversight, can be comprehensively resolved, thereby reducing the developers’ burden in understanding panic bugs and evaluating alternative functions.
5.3. Answer to RQ3: Efficiency
@@ 5.4. Answer to RQ4: Case Study
  • 1
    The full name of this crate is: unicode-segmentation.
@@
Bug Case1: Listing 9 shows the resolution of an arithmetic overflow error, which is a proposed patch of a closed issue for Rust crate ratatui (ratatui:online, ). Through fault localization, PanicKiller succeeded in pinpointing a list of potential fault locations, prioritizing src/layout.rs:233:20 as the top candidate, which aligns exactly with the actual fix. Then, PanicKiller successfully generated 3 patches for this location. In this case, the fault was identified within a struct filed expression; however, a closer inspection through iterative analysis revealed that this match expression included a binary expression, making it suitable for applying the Mutate Binary Operator pattern. Finally, we calculate the similarity score between the generated patch and the official patch, obtaining a score of $0.95$, indicating the correctness.
{listing}
[h] {minted}[linenos=false, fontsize=]diff - x: self.x + margin.horizontal, + x: 0 + self.x.saturating_add(margin.horizontal), y: self.y + margin.vertical, A generated patch for the crate ratatui that fixes a panic bug caused by arithmetic overflow.
{listing}
[h] {minted}[linenos=false, fontsize=]diff - let mut cur_item = self.items[cur_item_idx].unwrap(); + let mut cur_item = match self.items[cur_item_idx] + Some(item) =¿ item, + None =¿ return Err(Errors::ParseError + (ParsingErrors::UnknownFormat)), + ; A generated patch that fixes a panic bug in hifitime.
Bug Case 2: The patch illustrated in Listing 9, generated by PanicKiller, serves as an example of addressing a panic bug that arises from unwrapping a None/invalid value. This bug comes from an open issue of Rust crate hifitime (hifitime:online, ), and our patch has been merged by developers. PanicKiller employs the Insert Match Unwrapper pattern, specifically designed to mitigate problems stemming from the misuse of unwrap. This approach transitions the unsafe usage of unwarp() to pattern matching, which results in returning an error message rather than triggering a panic error. Additionally, the patch created by PanicKiller takes into consideration the variable types to ensure consistency with the code context, thereby guaranteeing that the modified program passes regression testing.
@@ 6.1. Application Scenarios
  • Benchmark dataset for the Rust APR tools. Panic4R is designed to serve as a robust and reliable benchmark dataset specifically tailored for the evaluation of APR tools targeting Rust panic bugs. It includes macro switches and test case execution scripts, facilitating researchers to employ it as a benchmark dataset in their evaluations.
  • Extendable dataset for Rust program repair. The project structure and test scripts included in Panic4R are both unified and highly extensible. Researchers interested in Rust and program repair can leverage this foundational dataset to expand into other bug types or develop new repair patches.
  • Templates for pattern-based APR tools development. The fix patterns can serve as templates for pattern-based APR tools. Based on this foundation, researchers can concentrate on developing improved techniques for fault localization and pattern prioritization. In this paper, we develop PanicKiller as an application to prove the effectiveness of these patterns.
  • Dataset for fine-tuning large language models (LLMs). As LLMs rapidly evolve, using high-quality data for fine-tuning can enhance their domain-specific applications. By leveraging our carefully curated Panic4R, which includes natural language interpretations and code change examples, researchers could fine-tune LLMs to develop AI-driven techniques for panic bug repair.
@@
  • Providing informative suggestions for users. PanicKiller produces a mix of suggestions—code patches and textual interpretations—that are more informative than compiler raw output. Given the steep learning curve of programs like Rust, we believe providing enhanced feedback is essential.
@@
  • Helpful for fixing real-world opening panic bugs. Our experiment results have shown PanicKiller is much more efficient than manual fixes. Additionally, PanicKiller has successfully resolved 28 panic bugs in real Rust crates, demonstrating its practical effectiveness.
@@ 6.2. Comparison with Existing Code Patterns
Comparison with Rust-specific code patterns. To date, only a few fix patterns have been specifically developed for Rust programs. Rust-lancet (yang2024lancet, ) and Ruxanne (robati2024patterns, ) are the primary studies that have designed bug and fix code patterns uniquely for Rust. Unlike these studies, which focus on ownership-related or other common bug types such as missing attributes, our infrastructure addresses the most critical panic bugs in Rust. While bugs violating lifetime rules and other common issues usually prevent compilation, offering error messages and corrective suggestions, panic bugs manifest at runtime, causing abrupt program termination. These bugs, influenced by Rust’s distinct memory and process management, render existing fix patterns inadequate for addressing panic issues. This paper introduces the first infrastructure aimed at exploring Rust panic bugs, identifying their root causes and developing specific repair patterns, significantly enhancing the understanding and remediation of Rust programs.
@@ 6.3. Threats to Validity
One potential threat to validity concerns the representativeness of our collected dataset and mined patterns, since all code and patches are sourced from open-source crates. However, we consider these crates to be relatively complex and representative of large-scale Rust projects. Panic4R comprises the top 500 most downloaded crates, reflecting the actual usage frequency and activity levels of these programs. We have also modeled our data collection process after the Defects4J dataset to enhance the reliability of our data. This adherence to proven methodologies in dataset construction supports the validity of our research findings.
@@ 7. Related Work
Automated Program Repair. Automated Program Repair (APR) has witnessed significant advancements in recent years. Most of methods (liu2019avatar, ; koyuncu2019ifixr, ; xia2022alapha, ; wen2018capgen, ; chen2017jaid, ; chen2021jaid, ; saha2019herclues, ; wong2021varfix, ; liu2019kpar, ; liu2019tbar, ) are designed for Java programs, with evaluated on Defect4J (just2014defect4j, ), a dataset of real bugs from open source Java programs. A recent study (huang2023survey, ) divides non-learning-based APR into three categories: search-based (weimer2013ae, ; qi2013trpautorepair, ; wen2018capgen, ; sidiroglou2015cp, ; le2012genprog, ), constraint-based (pei2011autofix, ; pei2014autofix, ; wei2010autofix, ; chen2017jaid, ; chen2021jaid, ; afzal2021sosrepair, ; mechtaev2015directfix, ) and template-based (tan2015relifix, ; le2016hdrepair, ; saha2019herclues, ). VarFix (wong2021varfix, ), a search-based way of observing which combinations of edit operations pass the test. Constraint-based techniques like Nopol (xuan2017nopol, ) and SemFix (nguyen2013semfix, ) transform the repair process into a constraint solving problem, reducing the search space. kPAR (liu2019kpar, ) and AVATAR (liu2019avatar, ) generate fix patterns collected by manual extraction and static violation analysis respectively, used by iFixR (koyuncu2019ifixr, ). TBar (liu2019tbar, ) is proposed to assess the qualitative and quantitative diversity of previous repair templates. As for learning-based APR, AlaphaRepair (xia2022alapha, ) achieves state-of-the-art results on both Java and Python programs via zero-shot learning. VulRepair (fu2022vulrepair, ) highlights the advancement of NMT-based automated vulnerability repairs with pre-trained models.
@@
Different from existing APR tools that focus on Java and C programs (huang2023survey, ), we represents the first dedicated infrastructure targeted at Rust panic bugs, including a real-world dataset Panic4R and an APR tool PanicKiller designed for practical application. Several tools (yang2024lancet, ; robati2024patterns, ; deligiannis2023rustassistant, ) have been proposed to address Rust compilation bugs. However, they do not specifically target panic bugs that occur during runtime. Considering the significant impact of Defects4J and the steep learning curve associated with Rust, we believe Panic4R would serve as a foundational resource for Rust’s research.
@@
Rust Program Testing and Analysing. Due to Rust’s innovative safety mechanism, new challenges have been posed in its testing and analyzing. RustSmith (sharma2023rustsmith, ) employs random program generation to test the Rust compiler. As for RULF (jiang2022rulf, ) and SyRust (takashima2021syrust, ), they concentrate on testing Rust crates by generating API sequences. In the realm of Rust program analysis, RUPTA (li2024rupta, ) introduces a context-sensitive pointer analysis framework for Rust, successfully applied to construct call graphs. Additionally, through static analysis, tools like SafeDrop (cui2023safedrop, ) and Rudra (bae2021rudra, ) detect memory safety issues in large-scale Rust programs, contributing to enhanced program robustness. In contrast, RustCheck (xia2023rustcheck, ) employs dynamic analysis techniques to uncover memory safety vulnerabilities.
@@ 8. Conclusion
In this paper, we introduce an infrastructure designed to address Rust panic bugs. We construct the first Rust program’s fixing dataset, Panic4R, containing $102$ real-world panic bugs and their patches. Additionally, we conduct pattern mining to identify Rust-specific fixing patterns. We also introduce an APR tool, PanicKiller, which effectively localizes faults and generates patches, outperforming commercial LLM-based tools. Moreover, PanicKiller has successfully resolved 28 open issues related to panics, all of which have been confirmed and merged by developers.

### v2

@@ 5.3.2. Fixing Effectiveness
Table 6. The number of panic-eliminated/plausible/correct patches of different tools on Panic4R with perfect locations.
@@
Eliminated Plausible Correct Eliminated Plausible Correct Eliminated Plausible Correct
Top-1 GPT-4 6 9.8% 4 6.6% 4 6.6% 3 7.3% 3 7.3% 3 7.3% 9 8.8% 7 6.9% 7 6.9%
GPT4-multi 6 9.8% 4 6.6% 4 6.6% 3 7.3% 3 7.3% 3 7.3% 9 8.8% 7 6.9% 7 6.9%
PanicKiller 33 54.1% 24 39.3% 7 11.5% 10 24.4% 7 17.1% 4 9.8% 43 42.2% 31 30.4% 11 10.8%
Top-3 GPT-4 6 9.8% 5 8.2% 4 6.6% 3 7.3% 3 7.3% 3 7.3% 9 8.8% 8 7.8% 7 6.9%
GPT4-multi 6 9.8% 5 8.2% 4 6.6% 3 7.3% 3 7.3% 3 7.3% 9 8.8% 8 7.8% 7 6.9%
PanicKiller 36 59.0% 27 44.3% 15 24.6% 13 31.7% 11 26.8% 9 22.0% 49 48.0% 38 37.3% 24 23.5%
Top-5 GPT-4 6 9.8% 5 8.2% 4 6.6% 4 9.8% 4 9.8% 4 9.8% 10 9.8% 9 8.8% 8 7.8%
GPT4-multi 6 9.8% 5 8.2% 4 6.6% 4 9.8% 4 9.8% 4 9.8% 10 9.8% 9 8.8% 8 7.8%
PanicKiller 36 59.0% 27 44.3% 15 24.6% 13 31.7% 11 26.8% 9 22.0% 49 48.0% 38 37.3% 24 23.5%
Regarding to the panic-eliminated and plausible patches, we believe that, when coupled with the textual interpretations derived from our mined patterns, panic-eliminated patches can aid developers repair their program efficiently. This is because the failure of some patches to pass the regression tests may be attributed to additional logic after the patch is applied, which is beyond APR tool’s capabilities. Such panic-eliminated fixes, with minimal manual oversight, can be comprehensively resolved, thereby reducing the developers’ burden in understanding panic bugs and evaluating alternative functions.
5.3.3. Efficiency
@@ 5.3.4. Case Study
Bug Case1: Figure 7 (a) shows the resolution of an arithmetic overflow error, which is a proposed patch of a closed issue for Rust crate ratatui (rat, 2024). Through fault localization, PanicKiller succeeded in pinpointing a list of potential fault locations, prioritizing src/layout.rs:233:20 as the top candidate, which aligns exactly with the actual fix. Then, PanicKiller successfully generated 3 patches for this location. In this case, the fault was identified within a struct filed expression; however, a closer inspection through iterative analysis revealed that this match expression included a binary expression, making it suitable for applying the Mutate Binary Operator pattern. Finally, we calculate the similarity score between the generated patch and the official patch, obtaining a score of $0.95$, indicating the correctness.
Figure 7. Patches generated by PanicKiller have fixed real-world issues.
Bug Case 2: The patch illustrated in Figure 7 (b), generated by PanicKiller, serves as an example of addressing a panic bug that arises from unwrapping a None/invalid value. This bug comes from an open issue of Rust crate hifitime (hif, 2024), and our patch has been merged by developers. PanicKiller employs the Insert Match Unwrapper pattern, specifically designed to mitigate problems stemming from the misuse of unwrap. This approach transitions the unsafe usage of unwarp() to pattern matching, which results in returning an error message rather than triggering a panic error. Additionally, the patch created by PanicKiller takes into consideration the variable types to ensure consistency with the code context, thereby guaranteeing that the modified program passes regression testing.
@@ 6.1. Application Scenarios
(1) Benchmark dataset for the Rust APR tools. The widely-used benchmark Defects4J has been crucial and influential for APR research in Java (Liu et al., 2019c; Koyuncu et al., 2019; Liu et al., 2019b; Xia and Zhang, 2022; Wen et al., 2018; Chen et al., 2017, 2021; Wong et al., 2021). Given the current limitations in the infrastructure for Rust program research, the dataset we propose, Panic4R, is designed to serve as a robust and reliable benchmark for subsequent studies. It includes macro switches and test case execution scripts, facilitating researchers to employ it as a benchmark dataset in their evaluations.
Panic4R offers detailed information for each case, encompassing the bug-triggering scenario and the exact patch applied to address the specific bug. This setup facilitates easy reproduction of the bug. Thus, Panic4R enables further analysis to uncover underlying patterns and common root causes of panic bugs. This can lead to better insights into Rust’s common vulnerabilities and areas where APR tools can focus or improve, ultimately contributing to the robustness of Rust codebases and reducing the likelihood of runtime panics in production software. Moreover, it establishes a uniform evaluation metric, allowing for more objective and comparable assessments of APR tool effectiveness. Researchers and practitioners can measure improvements in patch accuracy, bug-fix efficiency, and detection of bug patterns across different tools, fostering a clearer understanding of APR advancements in the Rust ecosystem.
(2) Extendable dataset for Rust program repair. Rust, as an emerging language with significant advantages in memory safety, increasingly research and approaches designed for Rust have been proposed. In recent years, some of the existing tools (Deligiannis et al., 2023; Yang et al., 2024; Robati Shirzad and Lam, 2024) have been developed for fixing unique bugs in Rust programs. However, due to the less mature infrastructure of Rust compared to other mainstream languages, there has been a lack of structured and open-source datasets. This gap has made it challenging to standardize testing, repairing, and verification processes, and the collected datasets and code snippets are often not reusable for subsequent research.
In our proposed PanicFI, the project structure and test scripts included in Panic4R are both unified and highly extensible, including executable code snippets, test cases, and validation scripts. Future studies that aim to incorporate new panic bug repair datasets can easily extend Panic4R using a standardized format. Similarly, for other types of bug repair datasets, maintaining a similar repository structure allows for quick adaptation to standardized dataset formats, minimizing redundant and tedious work.
(3) Templates for pattern-based APR tools development. Pattern-based (also known as ”template-based”) APR tools was proved to be the most effective method compared to other approaches, such as search-based and constraint-based schemes (Huang et al., 2023). After fault localization, the pattern-based APR tool could target these defects to select the corresponding fix templates to generate candidate patches. The mined patterns in our proposed PanicFI, containing 34 sub-patterns addressing 9 types of panic root causes, can serve as templates for pattern-based APR tools. Based on this foundation, researchers can concentrate on developing improved techniques for fault localization and pattern prioritization. For example, multiple potential patterns may match for the identified fault locations, necessitating further research into efficient pattern selection and validation. Additionally, ensuring the correctness of program semantics following the application of these patterns merits deeper investigation.
In this paper, we demonstrate the effectiveness and practicality of these mined patterns through the implementation of PanicKiller. Future work could focus on developing more refined repair tools based on these patterns, aimed at enhancing the accuracy of program semantics and improving repair efficiency.
(4) Dataset for fine-tuning large language models (LLMs). In recent years, several research have employed deep learning models for automated program repair (Xia and Zhang, 2022; Fu et al., 2022; Chi et al., 2022), and their results have demonstrated the feasibility of using AI technologies to assist in fixing programs. However, the effectiveness of these approaches is often constrained by several factors, such as the scale of the datasets for model training, the risk of overfitting due to model fine-tuning, and the loss of information associated with tokenizing unstructured data like code and text (Zhang et al., 2023). With the rapid advancement of large-scale models, it is highly potential that program repair techniques based on these LLMs will be implemented. This development could significantly enhance the precision and efficiency of tradition learning-based automated repairs, mitigating current limitations and opening new avenues for research in software maintenance.
To support research in LLM-based program repair, high-quality datasets are essential (Gupta et al., 2017; Berabi et al., 2021). The datasets and mined patterns included in our proposed PanicFI, are particularly well-suited for use as training data for LLMs. Notably, the patterns we have identified are diverse, encompassing 34 sub-patterns and covering 9 different root causes of panic errors. Moreover, our open-sourced patterns feature various data structures, including source code, abstract AST structures, and corresponding textual descriptions. These rich sources of information are highly beneficial for the learning and evolution of LLMs, providing a robust foundation for developing more effective automated program repair technologies.
(5) Providing informative suggestions for users. Our observations during data collection reveal that Rust developers, particularly beginners, often experience a sense of concern towards panic bugs. Additionally, the complex stack traces provided by the compiler can be difficult to understand. If tools were available to assist programmers in comprehending the causes of panic bugs and offering relevant repair suggestions, it would significantly reduce the complexity of understanding the program.
@@
In our PanicFI, PanicKiller produces a mix of suggestions—code patches and textual interpretations—that are more informative than compiler raw output. Given the steep learning curve of programs like Rust, we believe providing enhanced feedback is beneficial for developers. Such tools could demystify the error handling process, enabling developers to more effectively address and rectify issues, thereby enhancing their confidence and proficiency in using Rust.
@@
(6) Helpful for fixing real-world opening panic bugs. The infrastructure we propose, PanicFI, is closely aligned with real-world application scenarios. The Panic4R is derived entirely from open-source Rust projects within the ecosystem, ensuring its relevance and applicability. The patterns we have mined originate from Rust’s official code implementations, further validating the authenticity and utility of our infrastructure. Moreover, PanicKiller is capable of handling real, large-scale Rust projects. Our experiment results have shown PanicKiller is much more efficient than manual fixes, and PanicKiller has successfully resolved 28 panic bugs in real Rust crates, demonstrating its practical effectiveness.
@@ 6.2. Comparison with Existing Code Patterns
Comparison with Rust-specific code patterns. To date, only a few fix patterns have been specifically developed for Rust programs. Rust-lancet (Yang et al., 2024) and Ruxanne (Robati Shirzad and Lam, 2024) are the primary studies that have designed bug and fix code patterns uniquely for Rust. Unlike these studies, which focus on ownership-related or other common bug types such as missing attributes, our infrastructure addresses the most critical panic bugs in Rust. While bugs violating lifetime rules and other common issues usually prevent compilation, offering error messages and corrective suggestions, panic bugs manifest at runtime, causing abrupt program termination. These bugs, influenced by Rust’s distinct memory and process management, render existing fix patterns inadequate for addressing panic issues. This paper introduces the first infrastructure aimed at exploring Rust panic bugs, identifying their root causes and developing specific repair patterns, significantly enhancing the understanding and remediation of Rust programs.
@@ 6.3. Threats to Validity
One of the potential threats to validity concerns the representativeness of our collected dataset, since all code and patches are sourced from open-source crates. However, we consider these crates to be relatively complex and representative of large-scale Rust projects. Panic4R comprises the top 500 most downloaded crates, reflecting the actual usage frequency and activity levels of these programs. We have also modeled our data collection process after the Defects4J dataset to enhance the reliability of our data. This adherence to proven methodologies in dataset construction supports the validity of our research findings.
Another potential threats lies in the incompleteness of the mined fixing patterns. In real-world scenarios, specific bug triggers may have unique or diverse repair strategies. To enhance the comprehensiveness of our mined patterns, we systematically explore and extract from Rust’s official implementation code, which is considered to contain the most standard repair strategies. Additional patterns that may emerge in the future can also be easily integrated into our released PanicFI, further improving the effectiveness of pattern-based repair tools.
@@ 7. Related Work
In this section, we compare our work with other APR approaches, as well as the testing and analysing work for Rust program.
@@
Automated Program Repair. Automated Program Repair (APR) has witnessed significant advancements in recent years. Most of methods (Liu et al., 2019c; Koyuncu et al., 2019; Xia and Zhang, 2022; Wen et al., 2018; Chen et al., 2017, 2021; Saha et al., 2019; Wong et al., 2021; Liu et al., 2019a, b) are designed for Java programs, with evaluated on Defect4J (Just et al., 2014), a dataset of real bugs from open source Java programs. A recent study (Huang et al., 2023) divides non-learning-based APR into three categories: search-based (Weimer et al., 2013; Qi et al., 2013; Wen et al., 2018; Sidiroglou-Douskos et al., 2015; Le Goues et al., 2012), constraint-based (Pei et al., 2011, 2014; Wei et al., 2010; Chen et al., 2017, 2021; Afzal et al., 2021; Mechtaev et al., 2015) and template-based (Tan and Roychoudhury, 2015; Le et al., 2016; Saha et al., 2019). VarFix (Wong et al., 2021), a search-based way of observing which combinations of edit operations pass the test. Constraint-based techniques like Nopol (Xuan et al., 2017) and SemFix (Nguyen et al., 2013) transform the repair process into a constraint solving problem, reducing the search space. kPAR (Liu et al., 2019a) and AVATAR (Liu et al., 2019c) generate fix patterns collected by manual extraction and static violation analysis respectively, used by iFixR (Koyuncu et al., 2019). TBar (Liu et al., 2019b) is proposed to assess the qualitative and quantitative diversity of previous repair templates. As for learning-based APR, AlaphaRepair (Xia and Zhang, 2022) achieves state-of-the-art results on both Java and Python programs via zero-shot learning. VulRepair (Fu et al., 2022) highlights the advancement of NMT-based automated vulnerability repairs with pre-trained models.
@@
Different from existing APR tools that focus on Java and C programs (Huang et al., 2023), we represent the first dedicated infrastructure PanicFI targeted at Rust panic bugs. Several tools (Yang et al., 2024; Robati Shirzad and Lam, 2024; Deligiannis et al., 2023) have been proposed to address Rust compilation bugs. However, they are not specifically designed to target panic bugs that occur during runtime. Considering the significant impact of Defects4J and the steep learning curve associated with Rust, we believe PanicFI would serve as a foundational resource for Rust’s research.
@@
Rust Program Testing and Analysing. Due to Rust’s innovative safety mechanism, new challenges have been posed in its testing and analyzing. RustSmith (Sharma et al., 2023) employs random program generation to test the Rust compiler. As for RULF (Jiang et al., 2022) and SyRust (Takashima et al., 2021), they concentrate on testing Rust crates by generating API sequences. In the realm of Rust program analysis, RUPTA (Li et al., 2024) introduces a context-sensitive pointer analysis framework for Rust, successfully applied to construct call graphs. Additionally, through static analysis, tools like SafeDrop (Cui et al., 2023) and Rudra (Bae et al., 2021) detect memory safety issues in large-scale Rust programs, contributing to enhanced program robustness. In contrast, RustCheck (Xia et al., 2023) employs dynamic analysis techniques to uncover memory safety vulnerabilities.
@@ 8. Conclusion
In this paper, we introduce an infrastructure PanicFI designed to support fixing panic bugs of real-world Rust programs. We construct the first Rust program’s fixing dataset, Panic4R, containing $102$ real-world panic bugs and their patches. Additionally, we conduct pattern mining based on Rust compiler source code to identify Rust-specific fixing patterns. We also introduce an APR tool, PanicKiller, which effectively localizes faults and generates patches, outperforming commercial LLM-based tools. Moreover, PanicKiller has successfully resolved 28 open issues related to panics, all of which have been confirmed and merged by developers.
Data Availability
The dataset, mined patterns, the source code of PanicKiller, and experiment results can be found at: https://sites.google.com/view/panickiller/home.